OpenStudy (anonymous):

approximately how wide is the country of slovakia? A. 300 miles B. 100 miles C. 500 miles D. 600 miles

OpenStudy (anonymous):

@texaschic101

OpenStudy (texaschic101):

hmmm..it has a total area of 49,036 sq kilo or 18,933 sq miles..that a little bit bigger then the state of Mississippi. Its length is 539 km. A = L * W 49036 = 539 * W 49036/539 = W 90.97 = W.....my guess is about 100 miles

OpenStudy (anonymous):

thank you

OpenStudy (texaschic101):

don't thank me yet...wait till you see if it is right first...lol

OpenStudy (anonymous):

it is 300 miles